OpenStudy (anonymous):

How do you find the reciprocal of 3.6??

OpenStudy (anonymous):

you can turn 3.6 into a fraction, and then switch the numerator and denominator: 3.6 = 3.6/1 = 36/10 the reciprocal is 10/36 = 5/18 (i just reduced the fraction by dividing the top and bottom by 2) you could ALSO say the reciprocal of: 3.6/1 is 1/3.6, but i assume you're supposed to come up with a nice fraction...

OpenStudy (anonymous):

brilliant.